Suppose the following are national accounting data for a given year for Malaysia (hypothetical data) Amount (Billions SAUD) Consumption of fixed capital Gross private domestic investment Government consumption expenditures Government investment expenditures Imports Exports Household consumption expenditure Net property income paid overseas 285 725 720 165 -550 625 3010 -35 4 (a) (b) (c) (d) (e) Calculate GDP and the countrys gross national expenditure using the expenditure approach: Derive the countrys gross national product (GNP); Derive the countrys net national product (NNP); Derive the countrys current account balance; Derive the countrys gross national savings.
Answer #1

A). Through Expenditure approach

GDP = C + I + G + NX + gross domestic fixed capital formation

Where, C is household consumption expenditure

I is net private domestic investment

G is government consumption and investment expenditure

NX is export - imports

So, GDP = 720 + 165 + 725 + 3010 + 625 - 550 + 285 = 4980

So GDP is $4980.

GNE = GDP - Net exports

GNE = 4980 - 75 = 4905

So GNE is $ 4905.

b). GNP = GDP + net income from abroad = 4980 - 35 = 4945

So GNP is $4945.

c). Net national product = GDP + net income from abroad

= 4695 - 35 = $4640.

Net national product is $4640.

d). Current account balance = exports - imports = 625 - 550 = 75.

So CAD is $75.

e). Gross national savings;

GNS = GNP - private consumption - Government consumption

= 4945 - 3010 - 720 = $1215.

So, national savings is $1215.